OpenStudy (anonymous):

Using Integral Calculus A ball is thrown downward from a window that is 80 m above the ground with an initial velocity of -67meter/second; (a) when does the ball strike on the ground, and (b) with what speed will it strike on the ground.

OpenStudy (anonymous):

please help me

OpenStudy (perl):

g is approximately 9.81 m/s^2

OpenStudy (anonymous):

what it means g^w in above?

OpenStudy (perl):

g stands for the acceleration due to gravity , which is a constant on the surface of earth

OpenStudy (perl):

On the surface of earth acceleration due to gravity is about 9.81 m/s^2 Choosing up is positive d^2y/dt^2 = -9.81 m/s^2 integrating we have dy/dt = -9.81t + C y ' (0) = -67 -67 = -9.81*0 + C -67 = C therefore dy/dt = -9.81*t - 67 integrating again we have y(t) = -9.81t^2/2 - 67t + C y(0)=87 87 = -9.81*(0)^2/2 - 67(0) + C 87 = C y(t) = -9.81*t^2/2 -67t + 87

OpenStudy (anonymous):

i can't get it what is the second velocity and the time does the ball strike on the ground

OpenStudy (anonymous):

i thought if we estimate the gravity it must be 9.82 m/s^2

OpenStudy (perl):

the second derivative of position (also known as acceleration) is constant -9.81

OpenStudy (perl):

you can use 9.82 if you want

OpenStudy (anonymous):

ahhh okay, thank you so much

OpenStudy (anonymous):

why is it 87 is C?

OpenStudy (anonymous):

i can't get it why c is equal to 87.

OpenStudy (perl):

oh that was a typo

OpenStudy (perl):

On the surface of earth acceleration due to gravity is about 9.81 m/s^2 Choosing up is positive d^2y/dt^2 = -9.81 m/s^2 integrating we have dy/dt = -9.81t + C y ' (0) = -67 -67 = -9.81*0 + C -67 = C therefore dy/dt = -9.81*t - 67 integrating again we have y(t) = -9.81t^2/2 - 67t + C y(0)=80 80 = -9.81*(0)^2/2 - 67(0) + C 80 = C y(t) = -9.81*t^2/2 -67t + 80

OpenStudy (anonymous):

thanks.